OpenStudy (anonymous):

A few weeks into the deadly SARS (Severe Acute Respiratory Syndrome) epidemic in 2003, the number of cases was increasing by about 4% each day.† On April 1, 2003 there were 1,804 cases. Find an exponential model that predicts the number A(t) of people infected t days after April 1, 2003. A(t) = 1804(1.04^t) Use your model to estimate how fast the epidemic was spreading on April 17, 2003. (Round your answer to the nearest whole number of new cases per day.

OpenStudy (anonymous):

just plug in the number of days that had past to get your answer

OpenStudy (anonymous):

I did and I got 3,378.85 but its says to round to the nearest whole number and therefore its 3,379 cases per day, but my homework is online and its says its wrong

OpenStudy (anonymous):

try this 1804(1.04^17)=3,517.8

OpenStudy (anonymous):

i mean this 3518

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

the nearest whole number means no decimal

OpenStudy (anonymous):

ok try 3379

OpenStudy (anonymous):

oh nm

OpenStudy (anonymous):

yep i put 3518 and 3379 and still wrong

OpenStudy (anonymous):

3379 doesnt work either

OpenStudy (anonymous):

oh i think i got it. 3514

OpenStudy (anonymous):

nope hahaha

OpenStudy (anonymous):

@igreen we need help

OpenStudy (igreen):

What does 't' represent? Days?

OpenStudy (anonymous):

yes

OpenStudy (igreen):

\(A(17) = 1804(1.04^{17})\) \(A(17) = 1804(1.9479005)\) \(A(17) = 3514.0125\)

OpenStudy (igreen):

17 days is approximately $3,514.

OpenStudy (igreen):

I get that when I graph too..

OpenStudy (igreen):

3,514 or 3,513

OpenStudy (anonymous):

@iGreen nope he already tried 3514

OpenStudy (igreen):

Maybe $3,513 will work.

OpenStudy (anonymous):

hopefully.

OpenStudy (anonymous):

nope